# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
$X:=\left\{x\in V:x\in L_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}:=\left\{x\in(Y\cup Z)\setminus X:x\not\in Y_{i}\cup Z_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}^{\prime}:=\left\{x\in X_{YZ}:x\in W_{i}\leavevmode\nobreak\ \textrm{for some $W\in\{Y,Z\}$ with at least}\leavevmode\nobreak\ \frac{3}{2}\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
Recall that for all $i\in{\mathcal{C}}_{3}$, we have $|L_{i}|\leq 2{\epsilon}n$ and $|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|<4\delta n$. Hence
@@
$\leq$ $\sum_{i\in\mathcal{C}_{3}}\left(|(Y\cup Z)\setminus(Y_{1}\cup Z_{1})|+|(Y_{1}\cup Z_{1})\Delta(Y_{i}\cup Z_{i})|\right)$
@@
$X_{Y}:=$$\left\{x\in Y\setminus(X\cup X_{YZ}^{\prime}):x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 20\sqrt{\delta}|\mathcal{C}_{3}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}}_{3}\right\},$
$X_{Y}^{\prime}:=$$\left\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 6\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{3}}\right\}.$
By a similar discussion as above, we know that $|X_{Y}|<\frac{1}{2}\sqrt{\delta}n$ and each vertex of $X_{Y}\setminus X_{Y}^{\prime}$ lies in $Z_{i}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$. Furthermore, $|X_{Y}\setminus X_{Y}^{\prime}|\leq 4\delta n$. Then move vertices in $X_{Y}\setminus X_{Y}^{\prime}$ to $Z$. Hence $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n$.
Define $V_{\rm bad}:=X\cup(\bigcup_{(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}}X_{YZ}^{\prime})\cup(\bigcup_{k\in[4]}X_{C^{k}}^{\prime})$. Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+3\sqrt{\delta})n$. Moreover, each vertex in $C^{k}\setminus V_{\rm bad}$ lies in $C_{i}^{k}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in{\mathcal{C}}_{3}$.
The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path. The proof follows a similar argument to that of Claim 5.3, and thus we omit it.
@@
Let $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ be two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s}\in Y\setminus V_{\rm bad}$, $v_{1}\in Z\setminus V_{\rm bad}$, where $Y,Z\in\{C^{1},C^{2},C^{3},C^{4}\}$. Assume that $|\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)|>21\sqrt{\delta}n$ and $|C^{i}\setminus(V(P\cup Q)\cup V_{\rm bad})|>25\delta n$ if $w_{1}$ or $w_{1}^{\prime}$ is chosen from $C^{i}$ for some $i\in[4]$.
@@
    If $(Y,Z)\in\{(C^{1},C^{2}),(C^{3},C^{4})\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. Similar result holds for $(Y,Z)\in\{(C^{2},C^{3}),(C^{4},C^{1})\}$.
@@
    If $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(1,3),(1,4),(4,3)\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. Similar result holds for $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(3,1),(3,2),(2,1)\}$.
@@
    If $(Y,Z)=(C^{i},C^{i})$ with $i\in\{2,4\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and two vertices $w_{1}\in C^{i+2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
@@
    If $|N_{i}^{+}(x,C^{1}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ with center in $x$ and both endpoints in $C^{1}$. Move such $x$ to $C^{1}$.
@@
    Notice that each vertex in $C^{1}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ in-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$. If $|N_{i}^{+}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{4}$ with center in $x$ and endpoints in $C^{4},C^{2}$ respectively. Move such $x$ to $C^{1}$.
    Suppose (a)-(b) do not hold. Then for at least $(1-400\delta)|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|C^{3}\cup C^{4}|-400\delta n-|V_{\rm bad}|$ out-neighbors in $(C^{3}\cup C^{4})\setminus V_{\rm bad}$. Notice that each vertex in $C^{3}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ out-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
@@
    If $|N_{i}^{-}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed path $P_{3}$ or $P_{4}$ with center in $x$ and endpoints in $C^{2},C^{4}$ respectively. Move such $x$ to $C^{3}$. Similar result holds when $|N_{i}^{-}(x,C^{3}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, move such $x$ to $C^{3}$.
@@
    Suppose (a)-(c) do not hold. Then in each $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|(C^{1}\cup C^{4})\setminus V_{\rm bad}|-400\delta n-|V_{\rm bad}|$ in-neighbors inside $(C^{1}\cup C^{4})\setminus V_{\rm bad}$ for all but at most $(1-400\delta)|\mathcal{C}_{3}|$ colors $i\in\mathcal{C}_{3}$. Move such $x$ to $C^{2}$.
Applying the above operation for $j\in\{2,3,4\}$ in turn. Denote by $X^{\prime}$ the subset of $X$ consisting of vertices from item (d). Hence, for vertices in $X\setminus X^{\prime}$, there exists a set of disjoint rainbow directed $P_{3}$ or $P_{4}$ with centers in $X\setminus X^{\prime}$, each having both endpoints in $C^{k}$ with $k\in\{1,3\}$, or one endpoint in $C^{2}$ and the other in $C^{4}$. Moreover, $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n+2\sqrt{\epsilon}n$.
By the definition of $V_{\rm bad}\setminus X$, for vertices in $(V_{\rm bad}\setminus X)\cap Y$, we can find a set of disjoint rainbow directed $P_{3}$ with centers in $(V_{\rm bad}\setminus X)\cap Y$ and endpoints in $Y\setminus V_{\rm bad}$ (if $Y\in\{C^{1},C^{3}\}$) or in $(C^{2}\cup C^{4})\setminus(Y\cup V_{\rm bad})$ (if $Y\in\{C^{2},C^{4}\}$). Denote the rainbow directed path with center $x$ obtained above by $P_{x}$ for each $x\in V_{\rm bad}\setminus X^{\prime}$.
@@
Denote $s_{1}:=|X_{C^{2}C^{4}}^{\prime}\cap C^{2}|$ and $s_{2}:=|X_{C^{2}}^{\prime}|$. If $|C^{2}|-|C^{4}|-\sigma\leq s_{1}+2s_{2}$, then by moving $s_{1}^{\prime}$ vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and $s_{2}^{\prime}$ vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$ with $s_{1}^{\prime}+2s_{2}^{\prime}=|C^{2}|-|C^{4}|-\sigma$, we can get our desired result. Hence, it suffices to consider $|C^{2}|-|C^{4}|-\sigma>s_{1}+2s_{2}$. We first move all vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and all vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$. Then $(V_{\rm bad}\setminus X)\cap C^{2}=\emptyset$. Choose a set of disjoint maximal rainbow directed paths in $\mathcal{D}[C^{2}]\cup\mathcal{D}[C^{2},C^{1}]$, denote such a set by $\{Q_{1},\ldots,Q_{t}\}$. Assume $\{Q_{1}^{\prime},\ldots,Q_{k}^{\prime}\}$ is a set of disjoint rainbow directed $P_{3}$ in $\mathcal{D}[C^{2},C^{3}]$ with centers in $C^{2}$ and endpoints in $C^{3}$.
@@
If $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$, then there must exist an integer $k^{\prime}\in[k]$ and a set of disjoint rainbow directed subpaths of $\{Q_{1},\ldots,Q_{t}\}$, say $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}}\}$, such that $|E(\tilde{Q}_{1})|+\cdots+|E(\tilde{Q}_{t^{\prime}})|+k^{\prime}=|C^{2}|-|C^{4}|-\sigma$. That is, $|C^{4}|-t^{\prime}=|C^{2}|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-k^{\prime}-\sigma$. If the end vertex of $\tilde{Q_{i}}$ with $i\in[t^{\prime}]$ (resp. the start vertex of $Q_{i}^{\prime}$ with $i\in[k^{\prime}]$) lies in $V_{\rm bad}\setminus X^{\prime}$, then extend it by a vertex not in $V_{\rm bad}$. Using Claim 5.15 to connect the resulting $\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}},Q_{1}^{\prime},\ldots,Q_{k^{\prime}}^{\prime}$ into a single rainbow path $P^{0}$ with endpoints in $C^{2}$ and $C^{4}$ respectively, whose length is at most $190\delta n$. Therefore, $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-\sigma$, as desired.
@@
In each of the above cases, the remaining vertices in $C^{2}$ and $C^{4}$ are balanced. Since $|E(P^{0})|\leq 190\delta n$, by avoiding vertices in $V(P^{0})$, one may choose a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{P_{x}:x\in V_{\rm bad}\setminus(X^{\prime}\cup P^{0})\}$. For colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{0})$, we select a maximal rainbow matching $M$ inside $\mathcal{D}^{\pm}[C^{2},C^{4}]\cup\mathcal{D}[C^{1}]\cup\mathcal{D}[C^{3}]\cup(\bigcup_{i\in[4]}\mathcal{D}[C^{i},C^{i+1}])$ ensuring it avoids vertices in $V(P^{0}\cup\mathbf{P}^{0})$. If there exists a color $c\in\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$, then $D_{c}^{\pm}[C^{1},C^{3}]$ is almost an undirected complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost undirected complete graphs. For colors in $\mathcal{C}\setminus({\rm col}(P^{0}\cup M))$,we can choose a set of disjoint rainbow directed paths, say $\mathbf{P}^{1}$, with centers in $C^{1}$ and endpoints inside $C^{3}$, ensuring that at most one color in $\mathcal{C}_{\rm bad}$ is remains unused.
@@
$\bullet$ $\sigma=0$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|$. We first assume that $\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})=\emptyset$. By applying Claim 5.5 inside the digraph collection $\{D_{i}[C^{1}]\cup D_{i}[C^{3}]:i\in\mathcal{C}_{3}\setminus\{i_{1},i_{2},i_{3},i_{4}\}\}$ and avoiding vertices and colors in $Q_{0}\cup Q_{1}\cup Q_{2}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$, where $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. Connecting $Q_{1},Q_{2}$ by using Claim 5.15 (iv), one gets a rainbow directed path $P=x_{1}\ldots y_{2}$ such that $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.

### v2

@@ 1. $(\dagger)$
$X:=\left\{x\in V:x\in L_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 3\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}:=\left\{x\in(Y\cup Z)\setminus X:x\not\in Y_{i}\cup Z_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 10\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
$X_{YZ}^{\prime}:=\left\{x\in X_{YZ}:x\in W_{i}\penalty 10000\ \textrm{for some $W\in\{Y,Z\}$ with at least}\penalty 10000\ \frac{3}{2}\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{3}}\right\}.$
Recall that for all $i\in{\mathcal{C}}_{3}$, we have $|L_{i}|\leq 2{\epsilon}n$ and $|(Y_{1}\cup Z_{1})\triangle(Y_{i}\cup Z_{i})|<4\delta n$. Hence
@@
$\leq$ $\sum_{i\in\mathcal{C}_{3}}\left(|(Y\cup Z)\setminus(Y_{1}\cup Z_{1})|+|(Y_{1}\cup Z_{1})\triangle(Y_{i}\cup Z_{i})|\right)$
@@
$X_{Y}:=$ $\left\{x\in Y\setminus(X\cup X_{YZ}^{\prime}):x\not\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 20\sqrt{\delta}|\mathcal{C}_{3}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}}_{3}\right\},$
$X_{Y}^{\prime}:=$ $\left\{x\in X_{Y}:x\in Y_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}n\penalty 10000\ \textrm{colors}\penalty 10000\ i\in{\mathcal{C}_{3}}\right\}.$
By a similar discussion as above, we know that $|X_{Y}|<\frac{1}{2}\sqrt{\delta}n$ and each vertex of $X_{Y}\setminus X_{Y}^{\prime}$ lies in $Z_{i}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in\mathcal{C}_{3}$. Furthermore, $|X_{Y}\setminus X_{Y}^{\prime}|\leq 4\delta n$. We then move vertices in $X_{Y}\setminus X_{Y}^{\prime}$ to $Z$. Hence $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n$.
Define $V_{\rm bad}:=X\cup\big(\bigcup_{(Y,Z)\in\{(C^{1},C^{3}),(C^{2},C^{4})\}}X_{YZ}^{\prime}\big)\cup\big(\bigcup_{k\in[4]}X_{C^{k}}^{\prime}\big)$. Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+3\sqrt{\delta})n$. Moreover, each vertex in $C^{k}\setminus V_{\rm bad}$ lies in $C_{i}^{k}$ for at least $(1-20\sqrt{\delta})|{\mathcal{C}}_{3}|$ colors $i\in{\mathcal{C}}_{3}$.
The subsequent result allows us to connect two disjoint short rainbow directed paths into a single short rainbow directed path. Its proof follows the same lines as that of Lemma 6.3, and is therefore omitted.
@@
Let $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ be two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s}\in Y\setminus V_{\rm bad}$, $v_{1}\in Z\setminus V_{\rm bad}$, where $Y,Z\in\{C^{1},C^{2},C^{3},C^{4}\}$. Assume that $|\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)|>21\sqrt{\delta}n$, and that for all $i\in[4]$, if $C^{i}$ is a candidate set for selecting $w_{1}$ or $w_{1}^{\prime}$, then $|C^{i}\setminus(V(P\cup Q)\cup V_{\rm bad})|>25\delta n$.
@@
    If $(Y,Z)\in\{(C^{1},C^{2}),(C^{3},C^{4})\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in Y\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. A similar result holds for $(Y,Z)\in\{(C^{2},C^{3}),(C^{4},C^{1})\}$.
@@
    If $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(1,3),(1,4),(4,3)\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$. A similar result holds for $(Y,Z)=(C^{i},C^{j})$ with $(i,j)\in\{(3,1),(3,2),(2,1)\}$.
@@
    If $(Y,Z)=(C^{i},C^{i})$ with $i\in\{2,4\}$, then there are two colors $c_{1},c_{2}\in\mathcal{C}_{3}\setminus{\rm col}(P\cup Q)$ and a vertex $w_{1}\in C^{i+2}\setminus(V(P\cup Q)\cup V_{\rm bad})$ such that $u_{1}Pu_{s}w_{1}v_{1}Qv_{t}$ is a rainbow directed path with colors ${\rm col}(P\cup Q)\cup\{c_{1},c_{2}\}$.
@@
    If $|N_{i}^{+}(x,C^{1}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed copy of $P_{3}$ centered at $x$ with both endpoints in $C^{1}$. We then move such $x$ to $C^{1}$.
@@
    Notice that each vertex in $C^{1}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ in-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$. If $|N_{i}^{+}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then there is a rainbow directed copy of $P_{4}$ centered at $x$ with endpoints in $C^{4},C^{2}$ respectively. We then move such $x$ to $C^{1}$.
    Suppose (a)-(b) do not hold. Then for at least $(1-400\delta)|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|C^{3}\cup C^{4}|-400\delta n-|V_{\rm bad}|$ out-neighbors in $(C^{3}\cup C^{4})\setminus V_{\rm bad}$. Observe that every vertex in $C^{3}\setminus V_{\rm bad}$ has at least $|C^{4}\setminus V_{\rm bad}|-16\delta n$ out-neighbors inside $C^{4}\setminus V_{\rm bad}$ for all but at most $20\sqrt{\delta}|\mathcal{C}_{3}|$ digraphs $D_{\ell}$ with ${\ell}\in\mathcal{C}_{3}$.
@@
    If $|N_{i}^{-}(x,C^{2}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$, then we can find a rainbow directed copy of $P_{3}$ or $P_{4}$ centered at $x$ with endpoints in $C^{2},C^{4}$ respectively; we then move such $x$ to $C^{3}$. The same conclusion holds when $|N_{i}^{-}(x,C^{3}\setminus V_{\rm bad})|\geq 200\delta n$ for at least $200\delta n$ colors $i\in\mathcal{C}_{3}$; in that case we also move $x$ to $C^{3}$.
@@
    Suppose (a)-(c) do not hold. Then in each $D_{\ell}$ with $\ell\in\mathcal{C}_{3}$, $x$ has at least $|(C^{1}\cup C^{4})\setminus V_{\rm bad}|-400\delta n-|V_{\rm bad}|$ in-neighbors inside $(C^{1}\cup C^{4})\setminus V_{\rm bad}$ for all but at most $(1-400\delta)|\mathcal{C}_{3}|$ colors $i\in\mathcal{C}_{3}$. We then move $x$ to $C^{2}$.
We perform the above operation successively for $j=2,3,4$. Let $X^{\prime}$ denote the subset of $X$ formed by the vertices described in item (d). Hence, for the vertices in $X\setminus X^{\prime}$, there exists a set of disjoint rainbow directed copies of $P_{3}$ or $P_{4}$ centered at those vertices, where each copy either has both endpoints in the same part $C^{k}$ with $k\in\{1,3\}$, or has one endpoint in $C^{2}$ and the other in $C^{4}$. Moreover, $\left||C^{2}|-|C^{4}|\right|\leq 24\delta n+2\sqrt{\epsilon}n$.
By the definition of $V_{\rm bad}\setminus X$, for vertices in $(V_{\rm bad}\setminus X)\cap Y$, we can find a set of disjoint rainbow directed copies of $P_{3}$ with centers in $(V_{\rm bad}\setminus X)\cap Y$ and endpoints in $Y\setminus V_{\rm bad}$ (if $Y\in\{C^{1},C^{3}\}$) or in $(C^{2}\cup C^{4})\setminus(Y\cup V_{\rm bad})$ (if $Y\in\{C^{2},C^{4}\}$). Denote the rainbow directed path with center $x$ obtained above by $P_{x}$ for each $x\in V_{\rm bad}\setminus X^{\prime}$.
@@
Denote $s_{1}:=|X_{C^{2}C^{4}}^{\prime}\cap C^{2}|$ and $s_{2}:=|X_{C^{2}}^{\prime}|$. If $|C^{2}|-|C^{4}|-\sigma\leq s_{1}+2s_{2}$, then by moving $s_{1}^{\prime}$ vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and $s_{2}^{\prime}$ vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$ with $s_{1}^{\prime}+2s_{2}^{\prime}=|C^{2}|-|C^{4}|-\sigma$, we can get our desired result. Hence, it suffices to consider $|C^{2}|-|C^{4}|-\sigma>s_{1}+2s_{2}$. We first move all vertices in $X_{C^{2}C^{4}}^{\prime}\cap C^{2}$ to $C^{1}\cup C^{3}$ and all vertices in $X_{C^{2}}^{\prime}$ to $C^{4}$. Then $(V_{\rm bad}\setminus X)\cap C^{2}=\emptyset$. Choose a set of disjoint maximal rainbow directed paths in $\mathcal{D}[C^{2}]\cup\mathcal{D}[C^{2},C^{1}]$, denote such a set by $\{Q_{1},\ldots,Q_{t}\}$. Assume that $\{Q_{1}^{\prime},\ldots,Q_{k}^{\prime}\}$ is a set of disjoint rainbow directed copies of $P_{3}$ in $\mathcal{D}[C^{2},C^{3}]$ with centers in $C^{2}$ and endpoints in $C^{3}$.
@@
On the one hand, we consider that $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$. Then there exist an integer $k^{\prime}\in[k]$ and a set of disjoint rainbow directed subpaths of $\{Q_{1},\ldots,Q_{t}\}$, say $\{\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}}\}$, such that $|E(\tilde{Q}_{1})|+\cdots+|E(\tilde{Q}_{t^{\prime}})|+k^{\prime}=|C^{2}|-|C^{4}|-\sigma$. That is, $|C^{4}|-t^{\prime}=|C^{2}|-(|V(Q_{1}^{\prime})|+\cdots+|V(Q_{t^{\prime}}^{\prime})|)-k^{\prime}-\sigma$. If the end vertex of $\tilde{Q_{i}}$ with $i\in[t^{\prime}]$ (resp. the start vertex of $Q_{i}^{\prime}$ with $i\in[k^{\prime}]$) lies in $V_{\rm bad}\setminus X^{\prime}$, then we extend it by a vertex not in $V_{\rm bad}$. Applying Claim 7.1, we then connect the modified paths $\tilde{Q}_{1},\ldots,\tilde{Q}_{t^{\prime}},Q_{1}^{\prime},\ldots,Q_{k^{\prime}}^{\prime}$ into a single rainbow path $P^{0}$ with endpoints in $C^{2}$ and $C^{4}$ respectively, whose length is at most $190\delta n$. Therefore, $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|-\sigma$, as desired.
@@
In each of the above cases, that is, either $|C^{2}|-|C^{4}|-\sigma\leq s_{1}+2s_{2}$, or $|C^{2}|-|C^{4}|-\sigma>s_{1}+2s_{2}$ and $|E(Q_{1})|+\cdots+|E(Q_{t})|+k\geq|C^{2}|-|C^{4}|-\sigma$, the number of the remaining vertices in $C^{2}$ and $C^{4}$ are balanced. Since $|E(P^{0})|\leq 190\delta n$, by avoiding vertices in $V(P^{0})$, one may choose a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{P_{x}:x\in V_{\rm bad}\setminus(X^{\prime}\cup P^{0})\}$. For colors in $\mathcal{C}_{\rm bad}\setminus{\rm col}(P^{0})$, we select a maximal rainbow matching $M$ inside $\mathcal{D}^{\pm}[C^{2},C^{4}]\cup\mathcal{D}[C^{1}]\cup\mathcal{D}[C^{3}]\cup(\bigcup_{i\in[4]}\mathcal{D}[C^{i},C^{i+1}])$ ensuring it avoids vertices in $V(P^{0}\cup\mathbf{P}^{0})$. Notice that if there exists a color $c\in{\mathcal{C}_{\rm bad}}\setminus({\rm col}(P^{0}\cup M))$, then $D_{c}^{\pm}[C^{1},C^{3}]$ is almost an undirected complete bipartite graph and $D_{c}[C^{2}],D_{c}[C^{4}]$ are almost undirected complete graphs. Consequently, for colors in ${\mathcal{C}_{\rm bad}}\setminus({\rm col}(P^{0}\cup M))$, we can choose a set of disjoint rainbow directed paths, say $\mathbf{P}^{1}$, with centers in $C^{1}$ and endpoints inside $C^{3}$, ensuring that at most one color in $\mathcal{C}_{\rm bad}$ remains unused.
@@
$\bullet$ Suppose that $\sigma=0$, i.e., $|C^{4}\setminus V(P^{0})|=|C^{2}\setminus V(P^{0})|$. We first assume that $\mathcal{C}_{\rm bad}\setminus{\rm col}(M\cup\mathbf{P}^{0})=\emptyset$. By applying Lemma 6.2 to the digraph collection $\{D_{i}[C^{1}]\cup D_{i}[C^{3}]:i\in\mathcal{C}_{3}\setminus\{i_{1},i_{2},i_{3},i_{4}\}\}$, while avoiding vertices and colors in $Q_{0}\cup Q_{1}\cup Q_{2}$, we obtain two disjoint rainbow directed paths $P^{1}=x_{1}\ldots y_{1}$ and $P^{2}=x_{2}\ldots y_{2}$. Here $y_{1}\in N_{i_{3}}^{-}(u_{1},C^{1})$, $x_{2}\in N_{i_{4}}^{+}(v_{2},C^{3})$, $x_{1}\in C_{c_{1}}^{1}\setminus V_{\rm bad}$ and $y_{2}\in C_{c_{2}}^{3}\setminus V_{\rm bad}$ for two unused colors $c_{1},c_{2}\in\mathcal{C}_{3}$. We then connect $Q_{1},Q_{2}$ via Claim 7.1 (iv) to form a rainbow directed path $P=x_{1}\ldots y_{2}$, which satisfies $V_{\rm bad}\setminus X^{\prime}\subseteq V(P)$, $\mathcal{C}_{\rm bad}\subseteq{\rm col}(P)$ and $|C^{4}\setminus V(P)|=|C^{2}\setminus V(P)|+1$.